What is the magnitude of the forces when F=3, theta = 10 degrees. In kN?

у 4 kN F 15° 0 IT 30° х 6 kN
Answer #1

Solution:

у 4 kN 4* cos15 F*cos10 F 150 e F*sin 10 4* sin15 30° x 6 kN

Taking resoltion of horizontal forces,

F_{x}=6+(3*sin10^{o})-(4*sin15^{o})

F_{x}=5.486kN

Taking resolution of vertical forces,

F_{y}=(3*cos10^{o})+(4*cos15^{o})

F_{y}=6.818kN

Therefore, Resultant force is,

F=\sqrt{F_{x}^{2}+F_{y}^{2}}

F=\sqrt{5.486^{2}+6.818^{2}}

F=8.75kN